OpenStudy (anonymous):

16. Given 2.91 moles of a gas at 12 degrees Celsius and at 104.25 kilopascals, what is the volume of the gas? (The ideal gas constant is 0.0821 L · atm/mol · K; 101.325 Kpa = 1 atm) Select one: a. 6.60 x 101 L b. 6.53 x 10-1 L c. 2.79 x 100 L d. 2.75 x 10-2 L

OpenStudy (jfraser):

notice the units of the gas constant have both ATM and K, but the info you're given is kPa and Celsius. Convert the pressure to atm and the temp to celsius, then plug into the IGL:\[P (atm)*V (L) = n (moles)*R (\frac{L*atm}{mol*K})*T(K)\] so\[V(L) = \frac{n*R*T}{P}\]